Operations Report – December 2nd

SOL: 5

Name of person filing report: Ashley Kowalski

Non-nominal systems: N/A

Notes on non-nominal systems: N/A


Spirit rover used: Yes
Hours: 201.9 (Start), 202.0 (End)
Beginning charge: 100%
Ending charge: 91%
Currently charging: No

Opportunity rover used: No
Hours: N/A
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Curiosity rover used: No
Hours: N/A
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Perseverance rover used: Yes
Hours: 250.3 (Start), 250.5 (End)
Beginning charge: 100%
Ending charge: 87%
Currently charging: No

General notes on rovers: No additional notes/comments.

Summary of Hab operations:
WATER USE: 45 gallons used today, 155 gallons used in total thus far
Water (static tank): 395 gallons
Static tank pipe heater (on or off): On
Static tank heater (On or off): On
Toilet tank emptied: No.

Summary of internet: No issues with the internet. Connection is good.

Summary of suits and radios:

The crew used the two-piece EVA suits on the EVAs today. Before the first EVA, our Executive Officer, Alli Taylor, first put on Suit #6 and everything was working fine, but before she left the habitat, she noticed that she was getting warm and her helmet started to fog up. It seems that the fan had a very low air flow despite turning up the air flow dial, so she switched suits and took Suit #4 instead.

During the second EVA, our Health and Safety Officer, Barbara Braun, used Suit #4, as well. By the end of the EVA, she was experiencing low air flow. She asked another crew member to turn up her air flow dial, but by the time she returned to the front air lock at the end of the EVA, there was no air flow at all.

From a Crew Engineer’s perspective, I suspect this may be a charging issue because upon inspection of the suits tonight, I noticed that the suits in question were not set to the “off” mode and hence were not charging. I believe this is what caused the lack of airflow, but I will monitor the situation and report back tomorrow.

Finally, the two walkie talkies mentioned previous reports (one from the upstairs set and one from the downstairs set) are still not working.

Summary of GreenHab operations: Today, the GreenHab Officer harvested some carrot sprouts for dinner. Additional information can be found in the GreenHab Report.
WATER USE: ~28 gallons (~5 gallons today)
Heater: On
Supplemental light: Off
Harvest: 28.3 grams (28.3 oz; converted using 1 oz = 28.4 g)

Summary of ScienceDome operations: The Crew Engineer performed mirror inspections using the Aerospace Corp microscope and took images of the abnormalities/anomalies observed.
Dual split: No heat or AC used. All systems off.

Summary of RAM operations: N/A

Summary of any observatory issues: No observatory issues.

Summary of health and safety issues: No health and safety issues.

Questions, concerns and requests to Mission Support: The Crew Engineer has not yet installed the furnace air filter. Please advise if this is something you would like us to do tonight/tomorrow. Crew on second EVA today did an external inspection of the habitat. Everything looked good.

Copyright © The Mars Society. All rights reserved. | Main Site